OpenStudy (anonymous):

how to find geometric mean? question is... insert two G.M between 2 and 16

OpenStudy (amoodarya):

a1=2 a2 a3=16 a3 a2 -- =--- =q=ratio a2 a1 put a1,a3 into it to find a2

OpenStudy (amoodarya):

16 a2 -- = -- now find a2=geometric mean a2 2

hartnn (hartnn):

that gives you 1 GM. to find 2 GMs, we do somthing like this : \(\Large \dfrac{2}{x}=\dfrac{x}{y}=\dfrac{y}{16}\) find x and y which are your GMs

hartnn (hartnn):

doubts ?